OpenStudy (anonymous):

-2(1-5p)=or less 13?

OpenStudy (stamp):

- 2 ( 1 - 5p ) <= 13 Distribute. - 2 + 10p <= 13 Solve. 10p <= 15 p <= 15/10 Simplify. p <= 3/2